# Sharp Bounds for Multiple Models in Matrix Completion

### v2

@@ 4.4 Proof of Theorem 5
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 160\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
$+c_{3}\Bigg{(}40c_{3}\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+\frac{4\mu a^{2}\log d}{n}\Bigg{)},$
If $n\geq 160m\log^{4}d$, we have $\frac{4\sigma^{2}}{\tau^{2}}\leq\frac{1}{10}$ and $\frac{64a^{2}}{\tau^{2}}\leq\frac{2}{5}$. Then $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\frac{4\sigma^{2}}{\tau^{2}}\right)\geq\frac{4}{5}\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}$. Then we obtain that there is a constant $c_{4}$ such that the following holds with probability at least $1-\frac{1}{d}$,
$\frac{3}{5}\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq c_{4}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{5}s^{2}.$
This implies
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{3}s^{2}.$
Denote $\delta=\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$. We take $s^{2}=4\delta$. If $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}>s^{2}$, we can take $\tilde{A}_{H}=A_{0}+\eta(\hat{A}_{H}-A_{0})$ such that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}=\|\eta(\hat{A}_{H}-A_{0})\|_{w(F)}=s$. Note that $\tilde{A}_{H}$ also satisfies
$\|\tilde{A}_{H}-A_{0}\|_{1}\leq 3\sqrt{r}\|\tilde{A}_{H}-A_{0}\|_{F}.$
By Lemma F.2 in [Fan et al., 2018], we have
@@
$\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\delta+\frac{2}{3}s^{2}\leq\frac{11}{12}s^{2}.$
This contradicts with that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}=s^{2}$. Thus, we get
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{20c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$
@@
$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}\log d}{n}.$
4.5 Proof of Theorem 7 and 9
Based on [Klopp et al., 2017], (12) and (16) can be summarized in the same form
$(\hat{L},\hat{S})\in\underset{\begin{subarray}{c}\|L\|_{\infty}\leq\mathrm{a}\\ \|S\|_{\infty}\leq\mathrm{a}\end{subarray}}{\arg\min}\left\{\frac{1}{N}\sum_{i=1}^{N}\left(Y_{i}-\left\langle X_{i},L+S\right\rangle\right)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\mathcal{R}(S)\right\}.$
where $\mathcal{R}$ is certain matrix norm for $S$. Denote $\mathcal{R}^{*}(B)=\sup_{\mathcal{R}(A)\leq 1}\langle A,B\rangle$. [Klopp et al., 2017] offers the following general results for robust matrix completion.
Theorem 17.
Let Assumptions 1 and 4 be satisfied. Let $\lambda_{2}\geq 4\left(\mathcal{R}^{*}(\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i})+2a\mathcal{R}^{*}(\frac{1}{N}\sum_{i\in\Omega}X_{i})\right)$, and $\lambda_{1}>4\|\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}\|$. Then, with probability at least $1-\frac{4.5}{d}$,
$\frac{\left\|\hat{L}-L_{0}\right\|_{2}^{2}}{m_{1}m_{2}}+\frac{\left\|\hat{S}-S_{0}\right\|_{2}^{2}}{m_{1}m_{2}}\leq C\left\{\Psi_{1}+\Psi_{2}+\Psi_{3}\right\},$
where $C$ is an absolute constant, and $\Psi_{1}$ is defined as
$\Psi_{1}=\mu^{2}m_{1}m_{2}r\left(\alpha^{2}\lambda_{1}^{2}+\mathbf{a}^{2}\left(E\left\|\frac{1}{n}\sum_{i\in\Omega}\epsilon_{i}X_{i}\right\|\right)^{2}\right)+\mathbf{a}^{2}\mu\sqrt{\frac{\log(d)}{n}}.$
Here, we do not need to care about the definitions of $\Psi_{2}$ and $\Psi_{3}$ since they are unrelated to $O(\frac{rM}{n})$ part in the final bounds. The term $\frac{rM\log d}{n}$ only results from $\Psi_{1}$. Note $\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}=\frac{1}{\alpha n}\sum_{i\in\Omega}\xi_{i}X_{i}$. By (22) and (24), we can obtain
$\left\|\dfrac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}\right\|\leq C\frac{\sigma}{\alpha}\sqrt{\dfrac{1}{nm}}=C\sigma\sqrt{\frac{1}{\alpha Nm}}$
with probability at least $1-\frac{1}{d}$ and
$E\left\|\dfrac{1}{n}\sum_{i\in\Omega}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\dfrac{1}{nm}}.$
Plug the above two bounds in $\Psi_{1}$, then we remove the $\log d$ factor.
4.6 Proof of restricted strong convexity for $L^{2}$ loss function
In this section, $c_{0}\sim c_{4}$ are positive constants whose values do not change at each occurrence.
Lemma 18.
Recall $\|A\|^{2}_{w(F)}=\sum_{i,j}a_{ij}^{2}P_{ij}$. Assume $n\geq m\log^{6}d\log^{2}n$. For a non-zero matrix $A$, the following inequality holds, uniformly with probability at least $1-\dfrac{1}{d}$,
$\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}\geq\dfrac{3}{4}\|A\|_{w(F)}^{2}-c_{3}\left(\|A\|_{\infty}^{2}\dfrac{\mu\log d}{n}+\|A\|_{\infty}\|A\|_{1}E\|R\|\right),$ (32)
where $c_{3}$ is a universal positive constant and $R$ is defined as
$R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}.$
Proof of Lemma 18.
At first, we introduce a concentration inequality for empirical processes.
Lemma 19 (Theorem 3.27 in [Wainwright, 2019]).
$Z=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}f(Z_{i}).$
Without loss of generality, we can assume $\|A\|_{w(F)}=1$. For a pair of fixed parameters $(\alpha,\rho)$, define:
$S(\alpha,\rho)=\{A\in\mathbb{R}^{m_{1}\times m_{2}};\|A\|_{w(F)}=1,\|A\|_{\infty}\leq\alpha,\|A\|_{1}\leq\rho\}.$
Introducing the convenient short hand notation $F_{A}(X)=\langle X,A\rangle^{2}$, we define
$Z(\alpha,\rho)=\sup_{A\in S(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i})-\|A\|_{w(F)}^{2}\right|.$
For all $A\in S(\alpha,\rho)$, $|F_{A}(X_{i})|\leq\alpha^{2}$ for all $1\leq i\leq n$. Thus, we obtain
$\operatorname{Var}\left(F_{A}(X_{i})\right)$ $=E[F_{A}(X_{i})^{2}]-\left(EF_{A}(X_{i})\right)^{2}$
$\leq E[F_{A}(X_{i})^{2}]$
$\leq\alpha^{2}\|A\|^{2}_{w(F)}.$
Then we have $\tilde{\sigma}^{2}\leq\alpha^{2}$. Applying the concentration inequality (33) for the empirical processes with $t=\frac{2\mu\log d}{N}$, we get
$P\left(Z(\alpha,\rho)\geq c_{0}E[Z(\alpha,\rho)]+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+2c_{2}\frac{\alpha^{2}\mu\log d}{n}\right)$ $\leq\exp\left(-2\mu\log d\right).$
Now we estimate $E[Z(\alpha,\rho)]$. By Rademacher symmetrization and Ledoux-Talagrand contraction (see [Wainwright, 2019] Proposition 4.11) , we get
$E[Z(\alpha,\rho)]\leq 2E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle^{2}\Big{|}\leq 4\alpha E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle\Big{|}.$
Since $\|A\|_{1}\leq\rho$,
$E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle\Big{|}\leq\rho E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|=\rho E\|R\|,$
where we use $|\langle A,B\rangle|\leq\|A\|\|B\|_{1}$. Thus, we have
$P\left(Z(\alpha,\rho)\geq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+2c_{2}\frac{\mu\alpha^{2}\log d}{n}\right)\leq\frac{1}{d^{2\mu}}.$ (34)
Then we use the peeling argument to prove that the following:
$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\dfrac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|^{2}_{\infty}\log d}{n}$ (35)
holds uniformly with probability at least $1-\frac{1}{d}$. Let $\mathcal{E}$ be the event that the bound (35) is violated for some $A$ with $\|A\|_{w(F)}=1$. Note that under Assumption 2,
$\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}\|A\|_{w(F)}=\sqrt{\mu m_{1}m_{2}}.$
and
$\|A\|_{F}\geq\|A\|_{w(F)}=1.$
Then, we have
$\|A\|_{1}\leq\sqrt{m}\|A\|_{F}\leq\sqrt{\mu d^{3}}$
and
$1\leq\|A\|_{F}\leq\|A\|_{1}.$
Without loss of generality, we can assume $\|A\|\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
$\|A\|_{\infty}\leq\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}$
and
$\|A\|_{\infty}\geq\frac{\|A\|_{F}}{\sqrt{m_{1}m_{2}}}\geq\frac{1}{d}.$
For an integer pair $(k,l)$, define
$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{1}\leq 2^{l}\}$
Every matrix $A$ with $\|A\|_{w(F)}=1$ belongs to one $\tilde{S}(k,l)$ with $k=\lfloor\log_{2}\frac{1}{d}\rfloor,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d)\rceil$ and $l=1,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d^{3})\rceil$. Thus the total number of the pairs $(k,l)$ is less than $H^{2}=9\log_{2}^{2}(\mu d)$. Let $\mathcal{E}_{k,l}$ be the event that the bound (35) is violated for some $A\in\tilde{S}_{k,l}$. It can be seen that
$\mathcal{E}\subseteq\bigcup_{1\leq k\leq H\atop 1\leq l\leq H}\mathcal{E}_{kl}.$
Thus, if there exists a matrix $A$ with $\|A\|_{w(F)}=1$ that violates the bound (35), it must belong to some set $\tilde{S}_{k,l}$. Next, for $\rho=2^{l},\alpha=2^{k}$, define the event
~$\widetilde{\mathcal{E}}_{k,l}:=\left\{Z(\alpha,\rho)\geq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\dfrac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}\right\}.$
We claim that $\mathcal{E}_{k,l}\subseteq\widetilde{\mathcal{E}}_{k,l}$. Indeed, if event $\mathcal{E}_{k,l}$ occurs, then there must exist some $A\in\tilde{S}_{k,l}$ such that
$\left|\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\dfrac{\mu\log d}{n}}+16c_{2}\dfrac{\|A\|^{2}_{\infty}\mu\log d}{n}$
$\geq 16c_{0}2^{k-1}2^{l-1}E\|R\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\dfrac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
$\geq\tilde{4}c_{0}2^{k}2^{l}E\|R\|+\sqrt{2}c_{1}2^{k}\sqrt{\dfrac{\mu\log d}{n}}+4c_{2}2^{2k}\frac{\mu\log d}{n},$
showing that ~$\widetilde{\mathcal{E}}_{k,l}$ occurs. Putting together the pieces, we have, for $d$ sufficiently large,
$P(\mathcal{E})\stackrel{{\scriptstyle\text{ (i) }}}{{\leq}}\sum_{1\leq k\leq H\atop 1\leq l\leq H}P\left(\widetilde{\mathcal{E}}_{k,l}\right)\stackrel{{\scriptstyle\text{ (ii) }}}{{\leq}}\frac{H^{2}}{d^{2\mu}}\leq\frac{9\log^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\dfrac{1}{d},$
where inequality (i) follows from the union bound applied to the inclusion $\mathcal{E}\subseteq\bigcup_{k,\ell=1}^{H}\widetilde{\mathcal{E}}_{k,l}$; inequality (ii) is a consequence of the tail bound (34). So far, we have proved (35).
In the above argument, we normalize the matrix by its $\|\cdot\|_{w(F)}$ norm. Namely we consider $\frac{A}{\|A\|_{w(F)}}$, whose infinity norm and nuclear norm in (35) are $\dfrac{\|A\|_{\infty}}{\|A\|_{w(F)}}$ and $\dfrac{\|A\|_{1}}{\|A\|_{w(F)}}$. Multiplying $\|A\|_{w(F)}$ on both sides of formula (35), we obtain that with probability at least $1-\frac{1}{d}$,
$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|$ $\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}\sqrt{\dfrac{1}{nm}}+2\sqrt{2}c_{1}\|A\|_{\infty}\|A\|_{w(F)}\sqrt{\dfrac{\mu\log d}{n}}$
$+16c_{2}\|A\|^{2}_{\infty}\frac{\mu\log d}{n}.$
which implies
$\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}$ $\geq\|A\|_{w(F)}^{2}-16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|-\frac{\|A\|_{w(F)}^{2}}{4}-(8c_{1}^{2}+4c_{2})\|A\|^{2}_{\infty}\frac{\mu\log d}{n}$
$=\frac{3}{4}\|A\|_{w(F)}^{2}-16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|-(8c_{1}^{2}+4c_{2})\|A\|^{2}_{\infty}\frac{\mu\log d}{n}.$
@@ 4.7 Proof of restricted strong convexity for Huber loss function
In this section, $c_{0}\sim c_{3}$ are positive constants whose values do not change at each occurrence, $L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$.
@@
$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|+\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)},$ (36)
@@
$=\frac{1}{n}\sum_{i=1}^{n}(\psi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\psi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle$
$\geq\frac{1}{n}\sum_{i=1}^{n}(\psi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\psi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle\textbf{1}_{H_{i}}$
@@
$\geq G_{\frac{\tau}{4s}}\left(\left\langle X_{i},\frac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle\right)\textbf{1}_{\chi_{i}}.$
@@
$f(A)$ $\geq E[\langle X_{i},A\rangle^{2}]\textbf{1}\{|\langle X_{i},A\rangle|\leq\tau/4s\}\textbf{1}_{\chi_{i}}$
$\geq E\langle X_{i},A\rangle^{2}-E\langle X_{i},A\rangle^{2}\textbf{1}\{|\langle X_{i},A\rangle|>\tau/4s\}-E\langle X_{i},A\rangle^{2}\textbf{1}\{|\xi_{i}|>\tau/2\}$
@@
$\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)-f(A)\right|\leq 8c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (38)
for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ and $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
@@
$-\Bigg{(}8c_{0}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|+2\sqrt{2}c_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{w(F)}\sqrt{\frac{\mu\log d}{n}}$
@@
$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}+\frac{\mu\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}\log d}{n}\Bigg{)}-\frac{\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}}{10}$ (39)
@@
We follow a similar procedure as in the proof of Lemma 18. For fixed $(\alpha,\rho)\in\mathbb{R}^{2}$, define the space:
$S(\alpha,\rho)=\{A:\|A\|_{w(F)}^{2}=1,\|A\|_{1}\leq\rho,\|A\|_{\infty}\leq\alpha\}.$
@@
for $A\in S(\alpha,\rho)$. Recall $f(A)=E_{X_{i},\xi_{i}}[F_{A}(X_{i},\xi_{i})]$. Since $x^{2}\geq x^{2}\textbf{1}_{|x|\leq 2R}\geq G_{R}(x)$, for $A\in S(\alpha,\rho)$,
$\quad\leavevmode\nobreak\ E[F^{2}_{A}(X_{i},\xi_{i})]-f^{2}(A)$
$\leq E[F^{2}_{A}(X_{i},\xi_{i})]$
$\leq E[G^{2}_{\tau/4s}(\langle X_{i},A\rangle)]$
@@
$E[Z(\alpha,\rho)]$ $\stackrel{{\scriptstyle\text{i}}}{{\leq}}E\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}F_{A}(X_{i},\xi_{i})\right|$
$\stackrel{{\scriptstyle\text{ii}}}{{\leq}}2\alpha E\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i}\textbf{1}_{\chi_{i}},A\rangle\right|$
$\stackrel{{\scriptstyle\text{iii}}}{{\leq}}4\alpha\rho E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|.$
In step (i) we use symmetrization technique. By Property (ii) of $G_{R}(\cdot)$ and $\|A\|_{\infty}\leq\alpha$, $G_{\frac{\tau}{4s}}(\cdot)$ is Lipschitz with Lipschitz constant less than $2\alpha$. Then, in step (ii) we apply Talagrand-Ledoux contraction inequality. Step (iii) is due to $|\langle A,B\rangle|\leq\|A\|\|B\|_{1}$.

### v3

@@ 4.1 Proof of Theorem 1
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 160\sqrt{2}\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
$+c_{3}\Bigg{(}40c_{3}\mu a^{2}\mathbb{E}^{2}\|R\|rm_{1}m_{2}+\frac{4\mu a^{2}\log d}{n}\Bigg{)},$
If $n\geq 160m\log^{4}d$, we have $\frac{4\sigma^{2}}{\tau^{2}}\leq\frac{1}{10}$ and $\frac{64a^{2}}{\tau^{2}}\leq\frac{2}{5}$, which leads to $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\frac{4\sigma^{2}}{\tau^{2}}\right)\geq\frac{4}{5}\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}$. Then we obtain that there is a constant $c_{4}$ such that the following holds with probability at least $1-\frac{1}{d}$,
$\frac{3}{5}\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq c_{4}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{5}s^{2}.$
The above implies
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{3}s^{2}.$
Denote $\delta=\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$. We take $s^{2}=4\delta$. If $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}>s^{2}$, we can take $\tilde{A}_{H}=A_{0}+\eta(\hat{A}_{H}-A_{0})$ such that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}=\|\eta(\hat{A}_{H}-A_{0})\|_{w(F)}=s$. Note that $\tilde{A}_{H}$ also satisfies
$\|\tilde{A}_{H}-A_{0}\|_{*}\leq 3\sqrt{r}\|\tilde{A}_{H}-A_{0}\|_{F}.$
By Lemma F.2 in Fan et al. [2018], we have
@@
$\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\delta+\frac{2}{3}s^{2}\leq\frac{11}{12}s^{2},$
which contradicts with that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}=s^{2}$. Thus, we get
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{20c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$
@@
$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)rm_{1}m_{2}+\frac{\mu^{2}\log d}{n}.$
@@ 4.2 Proof of restricted strong convexity for Huber loss function
$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|R\|+\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)},$ (14)
@@
$=\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\phi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle$
$\geq\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\phi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle\textbf{1}_{H_{i}}$
@@
$\geq G_{\frac{\tau}{4s}}\left(\left\langle X_{i},\frac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle\right)\textbf{1}_{\chi_{i}},$
@@
$f(A)$ $\geq\mathbb{E}[\langle X_{i},A\rangle^{2}]\textbf{1}\{|\langle X_{i},A\rangle|\leq\tau/4s\}\textbf{1}_{\chi_{i}}$
$\geq\mathbb{E}\langle X_{i},A\rangle^{2}-\mathbb{E}\langle X_{i},A\rangle^{2}\textbf{1}\{|\langle X_{i},A\rangle|>\tau/4s\}-\mathbb{E}\langle X_{i},A\rangle^{2}\textbf{1}\{|\xi_{i}|>\tau/2\}$
@@
$\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}-f(A)\right|\leq 8c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (16)
for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ and $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
@@
$-\Bigg{(}8c_{0}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{w(F)}\sqrt{\frac{\mu\log d}{n}}$
@@
$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{*}\|\hat{A}_{H}-A_{0}\|_{\infty}+\frac{\mu\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}\log d}{n}\Bigg{)}-\frac{\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}}{10}$ (17)
@@
At first, we introduce a classical concentration inequality for empirical processes, which can be found in many textbooks such as Wainwright [2019].
Lemma 11 (Theorem 3.27 in Wainwright [2019]).
$Z=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}f(Z_{i}).$
For fixed $(\alpha,\rho)\in\mathbb{R}^{2}$, we define the space:
$S(\alpha,\rho)=\{A:\|A\|_{w(F)}^{2}=1,\|A\|_{*}\leq\rho,\|A\|_{\infty}\leq\alpha\}.$
@@
for $A\in S(\alpha,\rho)$. Recall $f(A)=\mathbb{E}_{X_{i},\xi_{i}}[F_{A}(X_{i},\xi_{i})]$. Since $x^{2}\geq x^{2}\textbf{1}_{|x|\leq 2R}\geq G_{R}(x)$, for $A\in S(\alpha,\rho)$,
$\quad\leavevmode\nobreak\ \mathbb{E}[F^{2}_{A}(X_{i},\xi_{i})]-f^{2}(A)$ $\leq\mathbb{E}[F^{2}_{A}(X_{i},\xi_{i})]$
$\leq\mathbb{E}[G^{2}_{\tau/4s}(\langle X_{i},A\rangle)]$
@@
$\mathbb{E}[Z(\alpha,\rho)]$ $\stackrel{{\scriptstyle\text{i}}}{{\leq}}\mathbb{E}\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}F_{A}(X_{i},\xi_{i})\right|$
$\stackrel{{\scriptstyle\text{ii}}}{{\leq}}2\alpha\mathbb{E}\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i}\textbf{1}_{\chi_{i}},A\rangle\right|$
$\stackrel{{\scriptstyle\text{iii}}}{{\leq}}4\alpha\rho\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|.$
In step (i) we use symmetrization technique. By Property (ii) of $G_{R}(\cdot)$ and $\|A\|_{\infty}\leq\alpha$, $G_{\frac{\tau}{4s}}(\cdot)$ is Lipschitz with Lipschitz constant less than $2\alpha$. Then, in step (ii) we apply Talagrand-Ledoux contraction inequality. Step (iii) is due to $|\langle A,B\rangle|\leq\|A\|\|B\|_{*}$.